### v2

@@ Abstract
The advancements in Generative Artificial Intelligence (GenAI) technologies such as ChatGPT provide opportunities to enrich educational experiences, but also raise concerns about academic integrity if misused. This study aims to explore how universities and educators respond and adapt to the development of GenAI in their academic contexts by analyzing academic policies and guidelines established by top-ranked US universities regarding the use of ChatGPT in higher education. The data sources include academic policies, statements, guidelines as well as relevant resources provided by the top 100 universities in the US. Results show that the majority of these universities adopt an open but cautious approach towards the integration of GenAI. Primary concerns lie in ethical usage, accuracy, and data privacy. Most universities actively respond and provide diverse types of resources, such as syllabus templates/samples, workshops, shared articles, and one-on-one consultations, with topics focusing on general technical introduction, ethical concerns, pedagogical applications, preventive strategies, data privacy, limitations, and detective tools. The findings provide two suggestions for educators in policy-making: establish discipline-specific policies, and manage sensitive information carefully, as well as four implications for educators in teaching practices: accept its presence, align its use with learning objectives, evolve curriculum to prevent misuse, and adopt multifaceted evaluation strategies rather than relying on AI detectors.
@@
Generative Artificial Intelligence, AI in education (AIED), Technology in education, Higher education, Academic integrity, Educational resources
@@ 1 Introduction
The development of artificial intelligence (AI) led to the rapid advancement of Large Language Models (LLMs) such as ChatGPT, GPT-4, Gemini, Claude 2, Llama 2, etc [2, 36, 37]. In November 2022, OpenAI first released ChatGPT [^1] https://openai.com/blog/chatgpt, which is a powerful language model-based chatbot that can understand human conversation and generate human-like texts. Since its release, ChatGPT has gained significant attention as well as vigorous discussion across a wide range of fields. In educational contexts, it can be employed through various applications, including generating ideas, revising grammatical errors, providing instant feedback, and evaluating and grading writing assignments [1, 9, 32]. However, the automatic generation of human-like texts also poses potential risks to academic integrity, especially when faced with writing-intensive assignments and language courses [27, 35]. Some educators express concerns about potential misuse by students, suggesting that students may rely heavily on ChatGPT, which might further impact their critical thinking and problem-solving abilities [17]. However, many other scholars and educators also view ChatGPT as an innovative teaching tool that can benefit writing instruction and learning [34]. Due to the nature of a new and emerging technology with constant changes and updates, teachers might find it difficult to effectively evaluate and integrate ChatGPT into their current teaching and learning environments. Thus, the question now forms from not only understanding what ChatGPT can do but also from what universities can offer and what faculty can apply in terms of guidance and strategies on the use of ChatGPT in educational academia. Specifically, it is necessary to examine how different universities and educators are currently perceiving, adapting to, and applying the use of such technology in higher education.
@@
This study aims to investigate what guidelines and resources are currently available for educators, teacher trainers, students, and researchers to adopt generative AI (GenAI) in their teaching, learning, and research in higher education. The findings will guide future policy-making, influence guideline development, as well as inform practitioners on how to integrate and manage GenAI in the rapidly evolving AI era.
@@ 2 Research questions
  • What are different universities’ perspectives regarding the use of ChatGPT and other GenAI tools?
  • What resources are currently available for both faculty and students to appropriately adopt ChatGPT and other GenAI tools in their classrooms?
3 Literature review
3.1 Impact of GenAI on Higher Education
Artificial intelligence (AI) and natural language processing (NLP) have experienced tremendous development in recent years. ChatGPT, a revolutionary technology created by OpenAI, is an advanced chatbot that uses AI and NLP techniques to generate coherent and human-like responses [16]. By using deep learning and neural networks, this technology is equipped to understand, analyze, and produce responses to a wide variety of prompts, including questions, statements, or academic inquiries, all within a few seconds. Due to its remarkable abilities, ChatGPT has emerged as a groundbreaking technology that rapidly attracted worldwide attention in various fields [16, 30, 31]. Undoubtedly, ChatGPT has been increasingly affecting higher education, as it has the potential to enhance learning experiences while also introducing some challenges to the current educational contexts [6, 10, 25].
3.2 Benefits of Integrating GenAI in Teaching and Learning
As GenAI shows its powerful capability in answering questions, auto-writing, and adapting to users’ needs [14], it brings plenty of potential benefits and opportunities for innovation to higher education. It encourages students to ask questions, clarify their needs, and delve into various topics as a self-regulated learning approach [39]. For students who prefer hands-on experiential learning, GenAI can serve as a useful learning resource that helps them independently solve problems and achieve goals [5]. In addition, GenAI can tailor its feedback based on individual student’s unique needs and specific learning contexts [29]. By analyzing the questions from students, GenAI can identify their learning contexts and adjust the answers accordingly. For example, when students need help with writing, GenAI can differentiate its feedback according to writing genres such as narrative and argumentation, offering genre-specific suggestions to meet individuals’ needs. Moreover, with appropriate instruction, GenAI can be a powerful technology to facilitate student engagement and collaboration in classrooms. For instance, it can generate different scenarios for students to work collaboratively to solve problems and further foster an interactive and supportive learning environment [32].
From the perspective of teachers, GenAI can also serve as an innovative teaching tool in their teaching practices. Depending on individual teacher’s needs and their own teaching contexts, GenAI can efficiently generate teaching materials, such as syllabi, lesson plans, homework prompts, and in-class activities [19, 28]. Therefore, teachers can save more time in teaching preparation and focus more on students’ performance. GenAI can also help teachers with assessment and feedback. For example, it can be used to generate grading rubrics that effectively address the learning objectives and concisely explain different grading levels. When it comes to homework and essay grading, GenAI can be trained to automatically grade students’ work. This feature was highlighted in a study by [18], where an advanced GenAI, such as ChatGPT, has demonstrated a high level of accuracy (correlation of 0.86 with human grading) in grading student essays. In terms of providing feedback, the AI system can further support teachers in providing personalized and timely feedback and tutoring that are based on each student’s individual learning needs and progress [3]. Overall, with appropriate introduction and training, GenAI has the potential to be a valuable resource for both teaching and learning in a higher education context.
3.3 Risks of Integrating GenAI in Teaching and Learning
Along with the possible benefits of using GenAI in higher education, the potential risks and ethical concerns should also be addressed. One of the primary concerns is how to ensure academic integrity, especially given that every student can easily access ChatGPT when they are writing essays and completing other assignments. Relying heavily on AI writing tools may pose potential risks of increased plagiarism and questions about authenticity. The misuse of GenAI for instantly generating essays and answers will lead to unethical academic practices and create inequities in educational contexts [22, 27, 29]. This issue becomes more complex for instructors. First, many students may mix AI-generated content with their own work, and some students may use the iterative process to guide AI to repeatedly refine the answers. Detecting such subtly AI-generated content requires a nuanced skill, which can be challenging for a lot of instructors. Second, some instructors may find it difficult to determine how much AI-generated content should be considered as plagiarism. What percentage of AI-generated content should be acceptable in students’ essays? What if students cite ChatGPT in their essays? Third, in order to avoid plagiarism and unethical academic practices, some instructors may struggle with designing assignments and evaluations that are less possible to use AI assistance, such as tasks that require deep critical thinking and need to meet some personalized requirements. However, instructors might need more time and effort in designing AI-resistant assignments, thus, it can create extra burdens and challenges for their daily teaching practices.
Another risk is that students’ over-reliance on GenAI might negatively impact their critical thinking and higher-order thinking skills. ChatGPT can always instantly respond to students’ requests and questions with personalized answers. When students get used to turning to ChatGPT for answers and essays, they might lose opportunities to engage deeply with the questions, and gradually lose abilities to develop their own analytical and problem-solving skills. Over-reliance on AI could also lead students to passively accept-generated answers without questioning and evaluation [9]. It could also reduce students’ motivation to brainstorm and create their own insights prior to answering questions or prompts. If students lack appropriate guidance from their instructors and universities, they might choose to turn to AI systems whenever they have the chance. Thus, this direction in their learning can eventually affect important learning abilities, such as problem-solving, decision-making, and creative thinking.
Furthermore, although GenAI can always produce quick responses, the answers may be superficial, biased, misleading, and even inaccurate. As a developing AI tool, the information processing system may still have limitations and can potentially lead to biased or incorrect outputs [11, 29]. When students use ChatGPT to generate quick answers, they may receive falsified information and develop an incorrect or superficial understanding. Without proper instruction, the misuse of ChatGPT may result in profound negative impacts on the overall quality of education and research.
Considering the threats and risks associated with GenAI, there is no doubt that there would be concerns and hesitations coming from researchers, educators, and policy-makers, especially when it comes to higher education contexts. Numerous studies have addressed the need for regulations and guidance to effectively integrate and manage the use of AI writing tools in schools and classrooms [9, 27, 32, 29, 15]. However, few studies have focused on reviewing and analyzing the current academic policies and resources regarding AI tools, and none of them have discussed the current resources offered by different universities. More research should be conducted to examine the policies and strategies that universities are implementing and to evaluate AI writing tools’ wider implications on higher education.
3.4 Academic Integrity Policies
Academic integrity, also known as honor codes, refers to the conduct of students, usually plagiarism and cheating, as well as the broader values, behavior, and practices of academics in all aspects of their work [20]. In the field of higher education, many universities often provide ethics education and training to pre-service and in-service teachers and faculty as a way to prepare them to teach. Universities and their faculty members often see academic integrity policies as a means to oversee and regulate student behaviors, rather than to examine or manage the values and conduct of the academic staff themselves. [20] suggest that academic integrity should be viewed as “a whole and entire” rather than being limited to student behavior. Specifically, it encompasses a wide range of aspects, including values, behaviors, and conduct of academics in teaching, research, and service.
Throughout history, academic integrity policies have been fundamental in fostering an academic community of honesty, fairness, truthfulness, and respect. With the release of ChatGPT and other AI technologies, some school administrators expressed concern that these tools may lead to a rise in academic dishonesty. They can serve as personal digital assistants that help students cheat during exams and plagiarize on other class assignments [35]. However, up to now, few studies have focused on exploring the academic integrity policies in the current technology era. Many universities have not established clear policies regarding the use of ChatGPT, but they are actively considering how to address this new technology [35]. This situation highlights the importance of an in-depth investigation of current existing policies and resources in different universities. This study examines academic policies and guidelines regarding AI writing tools, especially ChatGPT, from the top 100 prestigious universities in the US. The results will guide future policy development, teaching practices, and research in higher education.
@@ 4.1 Data Collection
The data in this study consists of policies, statements, resources, and guidelines regarding the use of GenAI, especially ChatGPT, from the top 100 U.S. universities listed in the 2024 US News Best National University Rankings. These universities represent a broad spectrum of reputational educational institutions across the U.S. To collect the data, we first identified the top 100 universities in the U.S. via the 2024 US News Best National University Rankings. Then we performed a systematic search using a list of keywords (e.g., GenAI policy, GenAI guidelines) together with the name of each university. The search results were filtered by the sources and content for each text. They were evaluated by the research team according to the inclusion and exclusion criteria (see Fig. 1). The inclusion criteria contain: 1) data from official university sources, such as the Office of Provost, Academic Senate, Center of Teaching and Learning, and Library Resources; 2) university-wide policies and statements regarding the use of GenAI tools and academic integrity; 3) guidelines and resources in relation to the use of ChatGPT and other GenAI tools in teaching, learning, and research. The exclusion criteria include: 1) articles from online news, and blog posts; 2) sources from a specific department or program of each university.
@@
Parent Codes
Child Codes
Definition
University Decision
Undecided/Unclear
The university has not made a clear decision or taken a definitive stance regarding GenAI.
Allow use with Conditions
The university permits the use of GenAI with conditions, such as appropriate citations.
Ban Use
The university prohibits the use of GenAI.
Instructor Decides
The university allows the use of GenAI depending on the instructor’s decisions.
Instructor Decision
Prohibition by Default
The use of GenAI is generally not allowed unless explicitly permitted by the instructor.
Permissibility by Default
The use of GenAI is generally allowed unless explicitly prohibited by the instructor.
Neutral
The university relies on the instructor’s decision without a specific stance.
Education Purpose
Plagiarism Prevention
To prevent students from directly copying texts generated from GenAI.
Authorship and Attribution
To require acknowledge AI-generated content in student academic assignments.
Limitations
To address limitations, including biased, inaccurate, unreliable, or falsely cited information generated by AI.
Research Purpose
Intellectual Property
To highlight the importance of acknowledging AI-generated content in professional research settings.
Data Privacy and Security
To address the confidentiality and security of data when using GenAI in professional research.
It is important to note that the policies and guidelines collected in this study are up until January 2024. We acknowledge that the development of GenAI policies and resources is dynamic and there will be changes regarding these policies in the future, however, as for our study, the current dataset serves as the foundation for our analysis. For those interested in exploring this dataset further, it is accessible at the dataset of university policies and resources.
@@ 4.2 Coding Schemes
Thematic analysis [4] was conducted to identify the themes related to universities’ perceptions and the availability of provided resources. Table 1 presents the coding scheme along with the definition for each code, designed for analyzing university policies and statements.
@@
Parent Codes
Child Codes
Definition
Target Audience
For Students
Resources provided specifically for students.
For Faculty
Resources provided specifically for faculty.
For General Audience
Resources provided specifically for the wider university community and public.
Types of Resources
Syllabus Templates and/or Examples
Suggested syllabus templates and/or examples shared on the resource and guideline pages.
Practical Training Workshop
Training workshops that train instructors and/or students to learn and try various functions of GenAI.
Dialogues and Discussions
Open dialogues and discussions for instructors and/or students to share their opinions.
Shared Articles and/or Blogs
Referenced articles and/or blogs that help instructors and/or students to further explore relevant topics.
One-on-one Consultations
Individual email, Zoom, and/or in-person consultations with school administration offices or representatives.
Content Analysis
General Technical Introduction
An overview focusing on the functions and technical aspects of GenAI.
Ethical Considerations
An introduction of ethical concerns on the use of GenAI.
Pedagogical Applications
Exploration of how GenAI can be incorporated into teaching and learning.
Preventive Strategies
Strategies to prevent students from using GenAI inappropriately.
Data Privacy
Guidelines for protecting instructors’ and students’ privacy when using GenAI for teaching and learning.
Limitations
Concerns on limitations, including biased, inaccurate, unreliable, or falsely cited information generated by AI.
Detective Tools
Introduction of available detective tools for detecting the use of GenAI.

### v3

@@ Abstract
The advancements in Generative Artificial Intelligence (GenAI) provide opportunities to enrich educational experiences, but also raise concerns about academic integrity. Many educators have expressed anxiety and hesitation in integrating GenAI in their teaching practices, and are in needs of recommendations and guidance from their institutions that can support them to incorporate GenAI in their classrooms effectively. In order to respond to higher educators’ needs, this study aims to explore how universities and educators respond and adapt to the development of GenAI in their academic contexts by analyzing academic policies and guidelines established by top-ranked U.S. universities regarding the use of GenAI, especially ChatGPT. Data sources include academic policies, statements, guidelines, and relevant resources provided by the top 100 universities in the U.S. Results show that the majority of these universities adopt an open but cautious approach towards GenAI. Primary concerns lie in ethical usage, accuracy, and data privacy. Most universities actively respond and provide diverse types of resources, such as syllabus templates, workshops, shared articles, and one-on-one consultations focusing on a range of topics: general technical introduction, ethical concerns, pedagogical applications, preventive strategies, data privacy, limitations, and detective tools. The findings provide four practical pedagogical implications for educators in teaching practices: accept its presence, align its use with learning objectives, evolve curriculum to prevent misuse, and adopt multifaceted evaluation strategies rather than relying on AI detectors. Two recommendations are suggested for educators in policy making: establish discipline-specific policies and guidelines, and manage sensitive information carefully.
@@
[…]
@@ 1 Introduction
The development of artificial intelligence (AI) led to the rapid advancement of large language models such as ChatGPT, GPT-4, Gemini, Claude 2, Llama 2, and etc (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023). In November 2022, OpenAI first released ChatGPT [^1] https://openai.com/blog/chatgpt, which is a powerful language model-based chatbot that can understand human conversation and generate human-like texts. Since its release, ChatGPT has gained significant attention as well as vigorous discussion across a wide range of fields. In educational contexts, it can be employed through various applications, including generating ideas, revising grammatical errors, providing instant feedback, and evaluating and grading writing assignments (Abdullayeva and Musayeva, 2023; Fuchs, 2023; Rudolph et al., 2023). However, the automatic generation of human-like texts also poses potential risks to academic integrity, especially when faced with writing-intensive assignments and language courses (Perkins, 2023; Sullivan et al., 2023). Some scholars express additional concerns about potential misuse by students, suggesting that students may rely heavily on ChatGPT, which might further impact their critical thinking and problem-solving abilities (Kasneci et al., 2023). Due to the nature of a new and emerging technology with constant changes and updates, many educators have also expressed anxiety and hesitation in integrating GenAI in their teaching practices, and are in need of recommendations and guidance from their universities that can support them to implement GenAI in classrooms effectively. Thus, the question now forms from not only understanding what ChatGPT can do but also from what universities can offer and what faculty can apply in terms of guidance and strategies on the use of ChatGPT in educational academia. Specifically, it is necessary to examine how different universities and educators are currently perceiving, adapting to, and applying the use of such technology in higher education.
@@
This study aims to investigate policies, guidelines, and resources currently provided by U.S. universities for their educators, teacher trainers, students, and researchers to adopt GenAI, especially ChatGPT, in their teaching, learning, and research in higher education. The findings will address numerous educators’ needs by informing AI-assisted teaching practices and guiding future policy-making, and thus impact the application of GenAI in higher education.
2 Literature review
Generative Artificial intelligence (GenAI) and natural language processing (NLP) have emerged as groundbreaking technologies that rapidly attract worldwide attention in various fields (Kalla and Smith, 2023; Ray, 2023; Rice et al., 2024). ChatGPT, a revolutionary technology created by OpenAI, is an advanced chatbot that uses AI and NLP techniques to generate coherent and human-like responses (Kalla and Smith, 2023). Using deep learning and neural networks, this technology is equipped to understand, analyze, and produce responses to a wide variety of prompts, including questions, statements, or academic inquiries, all within a few seconds.
GenAI has been increasingly affecting higher education, as it has the potential to enhance learning experiences while also posing challenges to the current educational contexts (Dempere et al., 2023; Grassini, 2023; Onal and Kulavuz-Onal, 2023). It encourages students to ask questions, clarify their needs, and delve into various topics as a self-regulated learning approach (Chiu, 2023; Cooper, 2023; Rasul et al., 2023; Wu et al., 2024). In the study conducted by Ng et al. (2024), students were taught to use ChatGPT at home to learn science concepts. The chatbot served as a resource to provide recommendations with science-related examples and explanations as well as scaffold the learning process by setting goals, suggesting learning strategies, and promoting time management skills (Ng et al., 2024). Another frequent pedagogical application of GenAI is as a writing assistant in teaching and learning academic writing (Crompton and Burke, 2023; Dempere et al., 2023; Imran and Almusharraf, 2023). They can assist student writers during the planning, drafting, and revising phases of academic writing, and provide suggestions that address their writing needs from linguistic nuances to genre-specific features (Alharbi, 2023; Liu et al., 2024; Mahapatra, 2024; Yan, 2023). As GenAI tools have been widely available for academics and students, concerns regarding academic integrity and copyright infringement become prominent in the discussions (Eke, 2023; Gao et al., 2023; Baek and Kim, 2023; Peres et al., 2023). Adopting AI-generated content without critical evaluation, rephrasing, and citation constitutes plagiarism and cannot be acceptable (Eke, 2023; Jarrah et al., 2023). In addition, some students experienced a lack of critical thinking and a decline in learning motivation due to overreliance on these tools (Escalante et al., 2023; Harunasari, 2022; Mahapatra, 2024; Song and Song, 2023). Therefore, the ongoing conversation highlights the misuse issues as a great concern that should be addressed. Institutions as well as teachers of higher education need to examine, adopt, and revise policies and strategies in their teaching contexts to preserve academic integrity, prevent plagiarism, and promote ethical implementation of GenAI tools.
While there is an increasing number of research experimenting with new technologies in education, there will always be anxiety and hesitation coming from teachers, especially due to the complexity and ethical issues of GenAI (Barrett and Pack, 2023). Studies by Iqbal et al. (2022) and Kiryakova and Angelova (2023) reported that a large portion of educators considered ChatGPT both a threat and a favorable opportunity and were cautious in using them. From the interviews in the study conducted by Iqbal et al. (2022), many teachers pointed out that there was a lack of training and support on the effective use of GenAI from their institutions, which was identified as a major obstacle to its application. Similarly, Barrett and Pack (2023) emphasized that deficiencies in GenAI policies and guidances from their institutions have led some instructors to hesitate to fully embrace GenAI due to ethical concerns and a naivety of how to appropriately use GenAI for educational purposes. Participating faculty in Van Wyk (2024) advocated that more resources, such as webinars, should be provided for faculty and students to raise GenAI awareness and literacy. Overall, teachers’ perspectives highlight the importance and urgency of examining available policies and resources on GenAI use in higher education, which will guide institutions in providing effective guidelines and preparing instructors for future curriculum design.
Standing in the midst of heated debates about the ethical use of GenAI in education, numerous studies are calling for further research on principles, strategies, and resources to harness GenAI as a value-driven opportunity to enhance learning (Ali et al., 2023; Firat, 2023; Lee et al., 2023; Michel-Villarreal et al., 2023; Van Wyk, 2024). However, limited research have focused on exploring how higher education institutions are currently perceiving, adapting to, and applying the use of GenAI. This study fills the research gap by exploring publicly available policies, statements, guidelines, and resources regarding GenAI provided by the top 100 universities across the U.S., and from there, discussing implications and suggestions for educators in policy-making and teaching practices.
@@ 3 Research questions
  • How do U.S. universities manage and regulate the use of ChatGPT and other GenAI tools in higher education contexts, and what perceptions can be found from their policies on integration?
@@
  • What resources and guidance do U.S. universities provide on the use of ChatGPT and other GenAI tools for teaching practices?
  • How do the school ranking tiers and academic specializations influence the trends in perceptions and resource provision regarding the use of GenAI in higher education and what pedagogical implications can we learn from them?
@@
Parent Codes Child Codes Definition
University Decision Undecided/Unclear The university has not made a clear decision or taken a definitive stance regarding GenAI.
Allow use with Conditions The university permits the use of GenAI with conditions, such as appropriate citations.
Ban Use The university prohibits the use of GenAI.
Instructor Decides The university allows the use of GenAI depending on the instructor’s decisions.
Instructor Decision Prohibition by Default The use of GenAI is generally not allowed unless explicitly permitted by the instructor.
Permissibility by Default The use of GenAI is generally allowed unless explicitly prohibited by the instructor.
Neutral The university relies on the instructor’s decision without a specific stance.
Education Purpose Plagiarism Prevention To prevent students from directly copying texts generated from GenAI.
Authorship and Attribution To require acknowledge AI-generated content in student academic assignments.
Limitations To address limitations, including biased, inaccurate, unreliable, or falsely cited information generated by AI.
Research Purpose Intellectual Property To highlight the importance of acknowledging AI-generated content in professional research settings.
Data Privacy and Security To address the confidentiality and security of data when using GenAI in professional research.
@@ 4.1 Data Collection
The data in this study consists of policies, statements, resources, and guidelines regarding the use of GenAI, especially ChatGPT, from the top 100 U.S. universities listed in the 2024 US News Best National University Rankings. These universities represent a broad spectrum of reputational educational institutions across the U.S. To collect the data, we first identified the top 100 universities (n = 104) in the U.S. via the 2024 US News Best National University Rankings. Then we performed a systematic search using a list of keywords (e.g., GenAI policy, GenAI guidelines) together with the name of each university. The search results were filtered by the sources and content for each text. They were evaluated by the research team according to the inclusion and exclusion criteria (see Fig. 1). The inclusion criteria contain: 1) data from official university sources, such as the Office of Provost, Academic Senate, Center of Teaching and Learning, and Library Resources; 2) university-wide policies and statements regarding the use of GenAI tools and academic integrity; 3) guidelines and resources in relation to the use of ChatGPT and other GenAI tools in teaching, learning, and research. The exclusion criteria include: 1) articles from online news, and blog posts; 2) sources from a specific department or program of each university.
@@
It is important to note that the policies and guidelines collected in this study are up until April 2024. We acknowledge that the development of GenAI policies and resources is dynamic and there will be changes regarding these policies in the future, however, as for our study, the current dataset serves as the foundation for our analysis. For those interested in exploring this dataset further, it is accessible at the dataset of university policies and resources.
@@ 4.2 Coding Schemes
Thematic analysis (Braun and Clarke, 2006) was conducted to identify the themes related to universities’ perceptions and the availability of provided resources. Table 1 presents the coding scheme along with the definition for each code, designed for analyzing university policies and statements.
@@
Parent Codes Child Codes Definition
Target Audience For Students Resources provided specifically for students.
For Faculty Resources provided specifically for faculty.
For General Audience Resources provided specifically for the wider university community and public.
Types of Resources Syllabus Templates and/or Examples Suggested syllabus templates and/or examples shared on the resource and guideline pages.
Practical Training Workshop Training workshops that train instructors and/or students to learn and try various functions of GenAI.
Dialogues and Discussions Open dialogues and discussions for instructors and/or students to share their opinions.
Shared Articles and/or Blogs Referenced articles and/or blogs that help instructors and/or students to further explore relevant topics.
One-on-one Consultations Individual email, Zoom, and/or in-person consultations with school administration offices or representatives.
Content Analysis General Technical Introduction An overview focusing on the functions and technical aspects of GenAI.
Ethical Considerations An introduction of ethical concerns on the use of GenAI.
Pedagogical Applications Exploration of how GenAI can be incorporated into teaching and learning.
Preventive Strategies Strategies to prevent students from using GenAI inappropriately.
Data Privacy Guidelines for protecting instructors’ and students’ privacy when using GenAI for teaching and learning.
Limitations Concerns on limitations, including biased, inaccurate, unreliable, or falsely cited information generated by AI.
Detective Tools Introduction of available detective tools for detecting the use of GenAI.